# Order isomorphisms of sup-stable function spaces: continuous, Lipschitz, c-convex, and beyond

### v2

@@ 1 Preliminaries
This result extends [AFG24, Theorem 5.1] originally written for anti-involutions only. The proof of 10 can be found in the Appendix on pProof of 10. 10 exclusively provides the existence of the anti-isomorphism and does not discuss its uniqueness, which in general does not hold. It is one of the goals of this article to show that in many cases the map $\bar{F}$ is nevertheless unique up to a translation $g$ and reparametrization $\phi$. Anti-involutions were described by [AAM09, Definition 11] as a “concept of duality”, this point of view being developed more at length in [AAM07]. 10 can also be seen as a functional analogue to [AASW23, Theorem 1.6] written for the power set $2^{\mathscr{X}}$. If there are no anti-involutions over a set $\mathcal{G}$, as that is the case for $\mathcal{G}$ the set of lower semicontinuous functions on a Hausdorff space with no isolated point, as shown in 32, then there is no such duality for $\mathcal{G}$. Describing anti-isomorphisms over complete lattices as an abstract nonlinear duality is actually at the core of abstract convexity theory as summarized in [Sin97]. In this article however, we do not use complete lattices since our sets are not inf-stable, unlike ¯$\overline{\mathbb{R}}^{\mathscr{X}}$, and we focus less on dualities and more on order isomorphisms.
@@ 2.1 The shape of order (anti)isomorphisms
We start with a simple result showing that, if we know one reference anti-isomorphism, as given for instance by 10, it is enough to study the order isomorphisms rather than the more arduous order anti-isomorphisms or anti-involutions.
@@
Let $A,B:\mathscr{F}\rightarrow\mathcal{G}$ be two order anti-isomorphisms over partially ordered sets $\mathscr{F}$ and $\mathcal{G}$, and set $J=A^{-1}B$. Then $J$ is an order isomorphism over $\mathscr{F}$, and we have that $B=AJ$ and $A=BJ^{-1}$. In particular, if there exists an anti-involution $\bar{F}:\mathcal{G}\rightarrow\mathcal{G}$, every anti-involution over $\mathcal{G}$ writes as $\bar{F}J$ with $J:\mathcal{G}\rightarrow\mathcal{G}$ an order isomorphism satisfying $\bar{F}J\bar{F}J=\operatorname*{Id}_{\mathcal{G}}$. The result also holds when “order” is replaced with (max,+).
@@
If $\mathcal{G}=\operatorname*{Rg}(B)$ as in (10), then, for every $f\in\operatorname*{Rg}(B)$, there exists $(w_{y})_{y\in\mathscr{Y}}$ such that $f(\cdot)=\sup_{y\in\mathscr{Y}}w_{y}+b(\cdot,y)$. Hence we have that $J(f)=\sup_{y\in\mathscr{Y}}J(w_{y}+b(\cdot,y))(\cdot)$ as $J$ commutes with the $\sup$, as shown in (9). Thus to characterize $J$, we just have to study its action on the functions $\{w+b(\cdot,y)\}_{y\in\mathscr{T},w\in\mathbb{R}_{\bot}}$. More generally this argument holds for sup-generators. However, unless there are other assumptions, such as continuity or compactness, see 26, the $\{w+b(\cdot,y)\}_{y\in\mathscr{T},w\in\mathbb{R}_{\bot}}$ are not necessarily sup-irreducible or relatively-inf-irreducible elements, which have the nice property of being stable under order isomorphisms.
@@
Every order isomorphism $J:\mathscr{F}\rightarrow\mathcal{G}$ over sup-stable sets ¯$\mathscr{F}\subset\overline{\mathbb{R}}^{\mathscr{Y}}$ and $\mathcal{G}\subset\overline{\mathbb{R}}^{\mathscr{X}}$ sends sup-irreducible elements (resp. $\mathscr{F}$-relatively-inf-irreducible) of $\mathscr{F}$ onto sup-irreducible elements (resp. $\mathcal{G}$-relatively-inf-irreducible) of $\mathcal{G}$.
@@
We just apply 7. Let $f_{0}\in\mathscr{F}$ be sup-irreducible in $\mathscr{F}$. Take $g_{1},h_{1}\in\mathcal{G}$ such that $Jf_{0}=\sup(g_{1},h_{1})$. Since $J$ is onto, we can find $g_{0},h_{0}\in\mathscr{F}$ such that $g_{1}=Jg_{0}$, $h_{1}=Jh_{0}$. Hence, using that $J$ commutes with the $\sup$, $Jf_{0}=J(\sup(g_{0},h_{0}))$. Since $J$ is bijective and $f_{0}$ is sup-irreducible, we have that $f_{0}=g_{0}$ or $f_{0}=h_{0}$, applying $J$, $Jf_{0}$ is thus sup-irreducible in $\mathcal{G}$. For $\mathscr{F}$-relatively-inf-irreducible functions $f_{0}$ we have similarly that if $Jf_{0}=\inf\nolimits^{\mathcal{G}}(g_{1},h_{1})$, then $f_{0}=\sup\{J^{-1}h\,|\,h\in\mathcal{G},\,h\leq g_{1},\,h\leq h_{1}\}$, so $f_{0}=\inf\nolimits^{\mathscr{F}}(J^{-1}g_{1},J^{-1}h_{1})$.
Concerning $T:\mathscr{F}\to\mathcal{G}$, if $Tf=\sup(g_{1},h_{1})$, then, by 7, $f=\inf\nolimits^{\mathscr{F}}(T^{-1}g_{1},T^{-1}h_{1})$, and we conclude as before. ∎
@@
In [AAFM12, Definitions 4.14,4.15], the authors say that a function $f\in\mathcal{G}$ is $P$-extremal if, for all $g,h\in\mathcal{G}$, $(f\leq g)\&(f\leq h)\Rightarrow g\leq h\text{ or }h\geq g$, and conversely $Q$-extremal if, for all $g,h\in\mathcal{G}$, $(f\geq g)\&(f\geq h)\Rightarrow g\leq h\text{ or }h\geq g$. In other words the upper/lower bounds on an extremal $f$ should be “comparable”. With the same proof as that of 13, one can show that $P$ and $Q$-extremal functions are preserved by order isomorphisms. Moreover, if $f$ is $P$-extremal, then $f=\inf\nolimits^{\mathcal{G}}(g,h)\Rightarrow f\leq g,h\stackrel{{\scriptstyle w.l.o.g}}{{\Rightarrow}}g\leq h\Rightarrow f=g$, so $f$ is relatively-inf-irreducible. For some sets, such as that of lower semicontinuous convex functions on a vector space, $P$-extremal functions and relatively-inf-irreducible ones coincide, since they are the “Dirac masses” $\delta^{\top}_{x}+\lambda$. However, for $1$-Lipschitz functions, there are simply no $P$-extremal functions, [^3] This is shown for instance by taking two points $x_{0}\neq x_{1}$. Then any $1$-Lipschitz function $f$ satisfies $f\leq\inf(d(\cdot,x_{0})+\lambda_{0},d(\cdot,x_{1})+\lambda_{0})$ with $\lambda_{0}=\max(f(x_{0}),f(x_{1}))$, but the two upper bounds are not comparable. whereas the distance functions $d(\cdot,x_{0})+\lambda$ are all relatively-inf-irreducible. This is why we favor the concept of irreducibility over comparability, even though the latter proved fruitful in [AAFM12, Section 4].
@@
Let $J:\mathscr{F}\rightarrow\mathcal{G}$ be an order isomorphism between some non-empty sets ¯$\mathscr{F}\subset\overline{\mathbb{R}}^{\mathscr{Y}}$ and $\mathcal{G}\subset\overline{\mathbb{R}}^{\mathscr{X}}$ stable by addition of scalars and finitely sup-stable. Assume also that
$\forall g\in\mathcal{G},\,\forall g^{\prime}=\lim_{\alpha}\uparrow g_{\alpha}\in\overline{\mathcal{G}}^{\sup}\text{ with $h_{\alpha}\in\mathcal{G}$, }(g\leq g^{\prime})\implies(\forall\epsilon>0,\,\exists\alpha_{\epsilon}\text{ s.t.\ }g\leq g_{\alpha_{\epsilon}}+\epsilon),$ (12)
@@
Then $J$ can be extended to an order isomorphism between ¯supsuperscript¯$\overline{\mathscr{F}}^{\sup}$ and $\overline{\mathcal{G}}^{\sup}$.
@@
For instance (12) is satisfied if we have that $\mathcal{G}$ is a subset of the continuous functions on a compact set $\mathscr{X}$. Indeed we then have that for any fixed $\epsilon>0$
$\forall x\in\mathscr{X},\exists\alpha_{x}\text{ s.t.\ }\epsilon+g_{\alpha_{x}}(x)\geq\min(g^{\prime}(x),g(x)),\text{ and }\\ \exists V_{x,\alpha_{x}}^{g}\text{ open neighborhood of $x$},\,\forall y\in V_{x,\alpha_{x}}^{g},\,\max(|g(x)-g(y)|,|g_{\alpha_{x}}(x)-g_{\alpha_{x}}(y)|)\leq\epsilon$
In other words: $\forall x\in\mathscr{X},\exists\alpha_{x},\,g\leq g_{\alpha_{x}}+3\epsilon$. Now since $\mathscr{X}\subset\bigcup_{x\in\mathscr{X}}V_{x,\alpha_{x}}^{g}$ and $\mathscr{X}$ is assumed compact, we can find a finite subcover indexed by $\{x_{1},\dots,x_{N}\}$ and $g\leq\sup_{i}g_{\alpha_{x_{i}}}+3\epsilon$ over $\mathscr{X}$. Since $g_{\alpha}$ was assumed to be increasing there is a $\bar{\alpha}$ such that $\sup_{i}g_{\alpha_{x_{i}}}\leq g_{\bar{\alpha}}$. Alternatively it can be seen as a consequence of Dini’s theorem applied to $g^{\prime}_{\alpha}=\sup(g,g_{\alpha})$. Eq.(12) still holds if $\mathscr{X}$ is no longer compact but the functions “tend to 00 at infinity”, i.e. for all $h\in\mathcal{G}$, $\epsilon>0$, there exists a compact $K_{\epsilon}\subset\mathscr{X}$ s.t. for $x\notin K_{\epsilon}$, $|h(x)|\leq\epsilon$.
@@
Let us show that $\tilde{f}\leq f$. Fix $h\in\mathscr{F}$ such that $J(h)\leq\sup_{h^{\prime}\leq f}J(h^{\prime})$. Set $g=J(h)$ and $g^{\prime}=\sup_{h^{\prime}\leq f}J(h^{\prime})$. Since $\mathcal{G}$ is finitely sup-stable, we can take $g^{\prime}=\lim\uparrow g_{\alpha}$ with each $g_{\alpha}$ a suprema of a finite number of functions $J(h^{\prime})$ satisfying $h^{\prime}\leq f$.color=blue!10,size=] PCA: détailler ? Hence by (12), for all $\epsilon>0$, there exists $g_{\alpha_{\epsilon}}$ such that $g\leq g_{\alpha_{\epsilon}}+\epsilon$, and $g_{\alpha_{\epsilon}}=\sup_{h_{i}^{\prime}\leq f,\,i\in\{1,\dots N\}}J(h_{i}^{\prime})$. Notice that in 7, (9) still holds in our setting for finite $\mathscr{A}$. Hence $J^{-1}$ commutes with a finite number of suprema. So we obtain $h\leq\sup_{i}J^{-1}(J(h^{\prime}_{i})+\epsilon)$. Applying (13), since $J(h^{\prime}_{i})\leq g^{\prime}$, for all $y\in\mathscr{Y}$, it holds $h(y)\leq\sup_{i}h^{\prime}_{i}(y)+v_{y}(\epsilon)\leq f(y)+v_{y}(\epsilon)$. Taking $\epsilon$ to 00, we get $h\leq f$.color=blue!10,size=] PCA: attention, si j’avais $\sup_{i}v_{y,i}(\epsilon)$, le nombre de $i$ dépendrait de $\epsilon$ ! C’est un peu pénible cette uniforme continuité.. Taking the supremum over $h$ shows that $\tilde{f}\leq f$.
To prove that $\tilde{f}\geq f$, we are going to show that, for any $f\in\overline{\mathscr{F}}^{\sup}$ and $g\in\overline{\mathcal{G}}^{\sup}$, we have that $g\geq\bar{J}f\implies\bar{J}^{1}g\geq f$. Take any $f,g$ satisfying $g\geq\bar{J}f$. Fix $f^{\prime}\in\mathscr{F}$ such that $f^{\prime}\leq f$, so $Jf^{\prime}\leq\bar{J}f\leq g$, whence $f^{\prime}\leq\bar{J}^{1}g$. Taking the supremum over $f^{\prime}$, we obtain $\bar{J}^{1}g\geq f$. Since $\bar{J}f\geq\bar{J}f$, the implication we have just shown entails that $\bar{J}^{1}\bar{J}f\geq f$. We thus obtained that $f=\bar{J}^{1}\bar{J}f$, so $\bar{J}$ is an order isomorphism.

### v3

@@ 1 Preliminaries
This result extends [AFG24, Theorem 5.1] originally written for anti-involutions only. The proof of ˜10 can be found in the Appendix on pProof of ˜10. ˜10 exclusively provides the existence of the anti-isomorphism and does not discuss its uniqueness, which in general does not hold. It is one of the goals of this article to show that in many cases the map $\bar{F}$ is nevertheless unique up to a translation $g$ and reparametrization $\phi$. Anti-involutions were described by [AAM09, Definition 11] as a “concept of duality”, this point of view being developed more at length in [AAM07]. ˜10 can also be seen as a functional analogue to [AASW23, Theorem 1.6] written for the power set $2^{\mathscr{X}}$. If there are no anti-involutions over a set $\mathcal{G}$, as that is the case for $\mathcal{G}$ the set of lower semicontinuous functions on a Hausdorff space with no isolated point, as shown in ˜31, then there is no such duality for $\mathcal{G}$. Describing anti-isomorphisms over complete lattices as an abstract nonlinear duality is actually at the core of abstract convexity theory as summarized in [Sin97]. In this article however, we do not use complete lattices since our sets are not inf-stable, unlike $\overline{\mathbb{R}}^{\mathscr{X}}$, and we focus less on dualities and more on order isomorphisms.
@@ 2.1 The shape of order (anti)isomorphisms
We start with a simple result showing that, if we know one reference anti-isomorphism, as given for instance by ˜10, it is enough to study the order isomorphisms rather than the more arduous order anti-isomorphisms or anti-involutions.
@@
Let $A,B:\mathscr{F}\rightarrow\mathcal{G}$ be two order anti-isomorphisms over partially ordered sets $\mathscr{F}$ and $\mathcal{G}$, and set $J=A^{-1}B$. Then $J$ is an order isomorphism over $\mathscr{F}$, and we have that $B=AJ$ and $A=BJ^{-1}$. In particular, if there exists an anti-involution $\bar{F}:\mathcal{G}\rightarrow\mathcal{G}$, every anti-involution over $\mathcal{G}$ reads as $\bar{F}J$ with $J:\mathcal{G}\rightarrow\mathcal{G}$ an order isomorphism satisfying $\bar{F}J\bar{F}J=\operatorname*{Id}_{\mathcal{G}}$. The result also holds when “order” is replaced with (max,+).
@@
If $\mathcal{G}=\operatorname*{Rg}(B)$ as in (10), then, for every $f\in\operatorname*{Rg}(B)$, there exists $(w_{y})_{y\in\mathscr{Y}}$ such that $f(\cdot)=\sup_{y\in\mathscr{Y}}w_{y}+b(\cdot,y)$. Hence we have that $J(f)=\sup_{y\in\mathscr{Y}}J(w_{y}+b(\cdot,y))(\cdot)$ as $J$ commutes with the $\sup$, as shown in (9). Thus to characterize $J$, we just have to study its action on the functions $\{w+b(\cdot,y)\}_{y\in\mathscr{Y},w\in\mathbb{R}_{\bot}}$. More generally this argument holds for sup-generators. However, unless there are other assumptions, such as continuity or compactness, see ˜26, the $\{w+b(\cdot,y)\}_{y\in\mathscr{Y},w\in\mathbb{R}_{\bot}}$ are not necessarily sup-irreducible or relatively-inf-irreducible elements, which have the nice property of being stable under order isomorphisms.
@@
Every order isomorphism $J:\mathscr{F}\rightarrow\mathcal{G}$ over sup-stable sets $\mathscr{F}\subset\overline{\mathbb{R}}^{\mathscr{Y}}$ and $\mathcal{G}\subset\overline{\mathbb{R}}^{\mathscr{X}}$ sends sup-irreducible elements (resp. $\mathscr{F}$-relatively-inf-irreducible) of $\mathscr{F}$ onto sup-irreducible elements (resp. $\mathcal{G}$-relatively-inf-irreducible) of $\mathcal{G}$.
@@
We just apply ˜7. Let $f_{0}\in\mathscr{F}$ be sup-irreducible in $\mathscr{F}$. Take $g_{1},h_{1}\in\mathcal{G}$ such that $Jf_{0}=\sup(g_{1},h_{1})$. Since $J$ is onto, we can find $g_{0},h_{0}\in\mathscr{F}$ such that $g_{1}=Jg_{0}$, $h_{1}=Jh_{0}$. Hence, using that $J$ commutes with the $\sup$, $Jf_{0}=J(\sup(g_{0},h_{0}))$. Since $J$ is bijective and $f_{0}$ is sup-irreducible, we have that $f_{0}=g_{0}$ or $f_{0}=h_{0}$, applying $J$, $Jf_{0}$ is thus sup-irreducible in $\mathcal{G}$. For $\mathscr{F}$-relatively-inf-irreducible functions $f_{0}$ we have similarly that if $Jf_{0}=\inf\nolimits^{\mathcal{G}}(g_{1},h_{1})$, then $f_{0}=\sup\{J^{-1}h\,|\,h\in\mathcal{G},\,h\leq g_{1},\,h\leq h_{1}\}$, so $f_{0}=\inf\nolimits^{\mathscr{F}}(J^{-1}g_{1},J^{-1}h_{1})$.
Concerning $T:\mathscr{F}\to\mathcal{G}$, if $Tf=\sup(g_{1},h_{1})$, then, by ˜7, $f=\inf\nolimits^{\mathscr{F}}(T^{-1}g_{1},T^{-1}h_{1})$, and we conclude as before. ∎
@@
In [AAFM12, Definitions 4.14,4.15], the authors say that a function $f\in\mathcal{G}$ is $P$-extremal if, for all $g,h\in\mathcal{G}$, $(f\leq g)\&(f\leq h)\Rightarrow g\leq h\text{ or }h\geq g$, and conversely $Q$-extremal if, for all $g,h\in\mathcal{G}$, $(f\geq g)\&(f\geq h)\Rightarrow g\leq h\text{ or }h\geq g$. In other words the upper/lower bounds on an extremal $f$ should be “comparable”. With the same proof as that of ˜13, one can show that $P$ and $Q$-extremal functions are preserved by order isomorphisms. Moreover, if $f$ is $P$-extremal, then $f=\inf\nolimits^{\mathcal{G}}(g,h)\Rightarrow f\leq g,h\stackrel{{\scriptstyle w.l.o.g}}{{\Rightarrow}}g\leq h\Rightarrow f=g$, so $f$ is relatively-inf-irreducible. For some sets, such as that of lower semicontinuous convex functions on a vector space, $P$-extremal functions and relatively-inf-irreducible ones coincide, since they are the “Dirac masses” $\delta^{\top}_{x}+\lambda$. However, for $1$-Lipschitz functions, there are simply no $P$-extremal functions, [^3] This is shown for instance by taking two points $x_{0}\neq x_{1}$. Then any $1$-Lipschitz function $f$ satisfies $f\leq\inf(d(\cdot,x_{0})+\lambda_{0},d(\cdot,x_{1})+\lambda_{0})$ with $\lambda_{0}=\max(f(x_{0}),f(x_{1}))$, but the two upper bounds are not comparable. whereas the distance functions $d(\cdot,x_{0})+\lambda$ are all relatively-inf-irreducible. This is why we favor the concept of irreducibility over comparability, even though the latter proved fruitful in [AAM11, Theorem 18] or [AAFM12, Section 4].
@@
Let $J:\mathscr{F}\rightarrow\mathcal{G}$ be an order isomorphism between some non-empty sets $\mathscr{F}\subset\overline{\mathbb{R}}^{\mathscr{Y}}$ and $\mathcal{G}\subset\overline{\mathbb{R}}^{\mathscr{X}}$ stable by addition of scalars and finitely sup-stable. Assume also that
$\forall g\in\mathcal{G},\,\forall g^{\prime}=\lim_{\alpha}\uparrow g_{\alpha}\in\overline{\mathcal{G}}^{\sup}\text{ with $g_{\alpha}\in\mathcal{G}$, }(g\leq g^{\prime})\implies(\forall\epsilon>0,\,\exists\alpha_{\epsilon}\text{ s.t.\ }g\leq g_{\alpha_{\epsilon}}+\epsilon),$ (12)
@@
Then $J$ can be extended to an order isomorphism between $\overline{\mathscr{F}}^{\sup}$ and $\overline{\mathcal{G}}^{\sup}$.
@@
Eq.(12) has to do with the regularity of the functions in $\mathcal{G}$, whereas (13) is a form of uniform continuity of $J$.
For instance, if $J^{-1}$ is $K$-Lipschitz for some $K>0$ in the sense $|J^{-1}(g+\epsilon)(y)-J^{-1}(g)(y)|\leq K\epsilon$ for all $\epsilon>0$, $g\in\mathcal{G}$, $y\in\mathscr{Y}$, then (13) holds. On the other hand, (12) is satisfied if we have that $\mathcal{G}$ is a subset of the continuous functions on a compact set $\mathscr{X}$, as a consequence of Dini’s theorem applied to $g^{\prime}_{\alpha}=\sup(g,g_{\alpha})$. Eq.(12) still holds if $\mathscr{X}$ is no longer compact but the functions “tend to 0 at infinity”, i.e. for all $h\in\mathcal{G}$, $\epsilon>0$, there exists a compact $K_{\epsilon}\subset\mathscr{X}$ s.t. for $x\notin K_{\epsilon}$, $|h(x)|\leq\epsilon$.
@@
Let us show that $\tilde{f}\leq f$. Fix $h\in\mathscr{F}$ such that $J(h)\leq\sup_{h^{\prime}\leq f}J(h^{\prime})$. Set $g=J(h)$ and $g^{\prime}=\sup_{h^{\prime}\leq f}J(h^{\prime})$. Since $\mathcal{G}$ is finitely sup-stable, we can take $g^{\prime}=\lim\uparrow g_{\alpha}$ with each $g_{\alpha}$ a supremum of a finite number of functions $J(h^{\prime})$ satisfying $h^{\prime}\leq f$. Hence by (12), for all $\epsilon>0$, there exists $g_{\alpha_{\epsilon}}$ such that $g\leq g_{\alpha_{\epsilon}}+\epsilon$, and $g_{\alpha_{\epsilon}}=\sup_{h_{i}^{\prime}\leq f,\,i\in\{1,\dots,N\}}J(h_{i}^{\prime})$. Notice that in ˜7, eq.(9) still holds in our setting for finite $\mathscr{A}$. Hence $J^{-1}$ commutes with a finite number of suprema. So we obtain
$h=J^{-1}(g)\leq J^{-1}(g_{\alpha_{\epsilon}}+\epsilon)=\sup_{i}J^{-1}(J(h^{\prime}_{i})+\epsilon).$
Applying (13), since $J(h^{\prime}_{i})\leq g^{\prime}$, for all $y\in\mathscr{Y}$, it holds $h(y)\leq\sup_{i}h^{\prime}_{i}(y)+v_{y}(\epsilon)\leq f(y)+v_{y}(\epsilon)$. Taking $\epsilon$ to 0, we get $h\leq f$. Taking the supremum over $h$ shows that $\tilde{f}\leq f$.
To prove that $\tilde{f}\geq f$, we are going to show that, for any $f\in\overline{\mathscr{F}}^{\sup}$ and $g\in\overline{\mathcal{G}}^{\sup}$, we have that $g\geq\bar{J}f\implies\bar{J}^{1}g\geq f$. Take any $f,g$ satisfying $g\geq\bar{J}f$. Fix $f^{\prime}\in\mathscr{F}$ such that $f^{\prime}\leq f$, so $Jf^{\prime}\leq\bar{J}f\leq g$, whence $f^{\prime}\leq\bar{J}^{1}g$. Taking the supremum over $f^{\prime}$, we obtain $\bar{J}^{1}g\geq f$ as claimed. Since $\bar{J}f\geq\bar{J}f$, the implication we have just shown gives that $\tilde{f}=\bar{J}^{1}\bar{J}f\geq f$. We thus obtained that $f=\bar{J}^{1}\bar{J}f$, so $\bar{J}$ is an order isomorphism.